SURVEYING CIVIL.

Please do by hand, no Excel, with explanation. Thank you.

Q#6 Prepare a set of level notes for the survey illustrated below Compute the elevations and provide the Arithmetic Check. (5 points) kod Rod Rod FS = 0.540 BS 1.977 FS= 0.982 BS-2.868 M 461 BM 460 Elev. 161.273 TP1 STA BS HI FS Elev.
Answer #1

elevation of BM 460 = 161.273, and BS of BM 460 = 2.868

so, height of instrument = RL of bench mark + BS of bench mark = 161.273+2.868 = 164.141

now, FS of TP1 = 0.982

so, elevation of TP1 = height of instrument - FS of TP1 = 164.141-0.982 = 163.159

Now, position of instrument is changed (1st Change Point),

so, height of instrument = elevation of TP1 + BS of TP1 = 163.159+1.977 = 165.136

so elevation of M461 = height of instrument - FS of M461 = 165.136-0.54 = 164.596

STA BS HI FS Elev
BM460 2.868 164.141 161.273
TP1 1.977 165.136 0.982 163.159
M461 0.540 164.596

Now, Arithmetic Check

summation of BS = 2.686 + 1.977 = 4.845

summation of FS = 0.982 + 0.540 = 1.522

so summation of BS - summation of FS = 4.845-1.522 = 3.323

also Last RL - 1st RL = 164.596 - 161.273 = 3.323

so here, so summation of BS - summation of FS = Last RL - 1st RL = 3.323

hence our calulation is correct.